1. let x1 x2 xn be a random sample from a population with pdf ... f(x theta) =theta. e power - theta .x

find cramer roa lower bound for variance of unbiased estimator of theta


1
Expert's answer
2021-02-24T06:04:00-0500

Given Probability density Function is-


"F(X,\theta)=\theta e ^{-\theta}"


Let "l(X,\theta)" be the natural logrithm of "f(x,\theta)"

"l=log(\theta e^{-\theta})=log\theta+loge^{-\theta}=log\theta-\theta"


Variance of unbiased estimator of \theta

"V(\hat{\theta})\ge \dfrac{1}{I(\theta)}~~~~~~~-(1)"


Where, the Fisher information "{\displaystyle I(\theta )}"  is defined by


"I(\theta)=-E[\dfrac{d^2l(X,\theta)}{d\theta^2}]"


Where, "E=" Expected value of samples


"I(\theta)=-E(-\dfrac{1}{\theta^2})"


"I(\theta)=\dfrac{E}{\theta^2}"


From eqs.(1) we have



"V(\hat{\theta})\ge \dfrac{1}{I(\theta)}"


"V(\hat{\theta})\ge\dfrac{\theta^2}{E}"